OpenStudy (anonymous):

If the cost function and demand curve for certain product are C(x) = 60x + 7200 and p = 300 – 2x respectively. Find Total revenue function

OpenStudy (ranga):

Isn't revenue the product of the cost function and demand curve?

OpenStudy (agent0smith):

^that's what i was thinking, but it's not based on any proof... just kinda made sense.

OpenStudy (anonymous):

the cost function doesn't affect the total revenue. TR=x*p(x).

OpenStudy (anonymous):

TR=price*quantity

OpenStudy (anonymous):

if you want to find the profit though, just subtract C(x) from TR.

OpenStudy (ranga):

Never took an economics class and I am trying to figure out what is what? Is x the number of units of the product? Is C(x) the cost to purchase or to produce?

OpenStudy (ranga):

And p(x) means?

OpenStudy (anonymous):

x is the number of units of the product. C(x) is the cost to produce.

OpenStudy (anonymous):

p(x) is the price charged for x units

OpenStudy (anonymous):

*for each of the x units being sold

OpenStudy (ranga):

Okay, in that case revenue = x * p(x) makes sense. So for this problem we don't really need C(x) since they don't ask for profit.

OpenStudy (asib1214):

revenue function????

OpenStudy (ranga):

Revenue function = x * p(x) = x * (300 – 2x) = 300x - 2x^2

OpenStudy (anonymous):

Thanks

OpenStudy (ranga):

I looked up the definition of a demand curve: it is a curve drawn with price per unit of the product along the y axis and the quantity demanded on the x axis. As the demand increases the price per product will decrease as is evidenced by the negative slope of 300 - 2x. So it makes sense to multiply x and p(x) to get the total revenue function.

OpenStudy (agent0smith):

That seems odd... doesn't high demand allow for higher price per unit? low demand should dictate low cost per unit

OpenStudy (ranga):

When things are produced in small numbers, there is a lot of overhead involved and the company has to charge a lot more per unit to cover the production cost and make some profit. As the demand increases, the price charged per unit will decrease. But overall they will make a lot more profit because they are selling a lot more. We can see that in real cases. Whenever a new product is introduced it costs a lot. Then as it becomes popular the price goes down.

OpenStudy (anonymous):

If the cost function and demand curve for certain product are C(x) = 60x + 7200 and p = 300 – 2x respectively. Find Total revenue function P(x)=R(x)-C(x) Where P(x)=profit, R(x)=Revenue (R(x)=xp, where p is the price C(x)=cost (a+m(x)), m(x)=margin cost P(x)=R(x)-C(x) P(x)=(x(300 – 2x))-(60x + 7200) P(x)=300x – 2x^2-60x -7200 Profit equation will be: P(x)=-2x^2+240x-7200 Revenue rquation R(x)=xp R(x)=x(300 – 2x) R(x)=300x-2x^2 Cost equation which you gave me. C(x)=60x + 7200

OpenStudy (ranga):

High demand and low supply then the price will be higher High demand and high supply the price will be lower.